# MDAR: A Multi-scene Dynamic Audio Reasoning Benchmark

### v1

@@ Abstract
The ability to reason from audio, including speech, paralinguistic cues, environmental sounds, and music, is essential for AI agents to interact effectively in real-world scenarios. Existing benchmarks mainly focus on static or single-scene settings and do not fully capture scenarios where multiple speakers, unfolding events, and heterogeneous audio sources interact. To address these challenges, we introduce MDAR, a benchmark for evaluating models on complex, multi-scene, and dynamically evolving audio reasoning tasks. MDAR comprises 3,000 carefully curated question–answer pairs linked to diverse audio clips, covering five categories of complex reasoning and spanning three question types. We benchmark 26 state-of-the-art audio language models on MDAR and observe that they exhibit limitations in complex reasoning tasks. On single-choice questions, Qwen2.5-Omni (open-source) achieves 76.67% accuracy, whereas GPT-4o Audio (closed-source) reaches 68.47%; however, GPT-4o Audio substantially outperforms Qwen2.5-Omni on the more challenging multiple-choice and open-ended tasks. Across all three question types, no model achieves 80% performance. These findings underscore the unique challenges posed by MDAR and its value as a benchmark for advancing audio reasoning research. Code and benchmark can be found at https://github.com/luckyerr/MDAR.
@@ 1 Introduction
Existing benchmarks mainly focus on static or single-scene settings and do not fully capture scenarios where multiple speakers, unfolding events, and heterogeneous audio sources interact. Prior work has advanced audio perception and reasoning research, but most existing datasets are restricted to single-scene scenarios, short audio clips, or narrowly scoped tasks (Sakshi et al., 2025; Ma et al., 2025b; Wang et al., 2025b). These limitations make it challenging to evaluate models on complex, temporally evolving audio reasoning, leaving a gap in understanding model capabilities in realistic multi-scene environments.
To address these challenges, we introduce MDAR, a benchmark for evaluating models on challenging, multi-scene, and dynamically evolving audio reasoning tasks. As shown in Figure 1, MDAR comprises 3,000 carefully curated question–answer pairs linked to diverse audio clips, covering five categories of complex reasoning in dynamic audio scenes: Scene Understanding, Social Relationships and Social Rreasoning, Event Reasoning, Temporal Reasoning, and Anomaly Detection and Safety. Each category targets specific aspects of cognitive reasoning, from inferring social roles and intentions to predicting causal and sequential events. In addition to the common single-choice questions, MDAR also includes open-ended questions that require free-form answers and multiple-choice questions with multiple audios for the first time.
Figure 1: Overview of MDAR benchmark. MDAR focuses on five types of complex, multi-scene, and dynamically evolving audio reasoning tasks, spanning three challenging question formats, to test the advanced reasoning, perception, and knowledge capabilities of existing audio models.
We benchmark 26 state-of-the-art audio language models on MDAR and observe that they exhibit limitations in complex reasoning tasks. In the single-choice setting, Qwen2.5-Omni (open-source) achieves 76.67% accuracy, whereas GPT-4o Audio (closed-source) reaches 68.47%. Among different tasks of single-choice questions, Qwen2.5-Omni (Xu et al., 2025a) performs worst on temporal reasoning (71.43%), while GPT-4o Audio struggles most on scene reasoning (61.27%). However, GPT-4o Audio substantially outperforms Qwen2.5-Omni on the more challenging multiple-choice and open-ended tasks. Across all three question types, no model surpasses 80% performance. Overall, these results indicate that even the strongest available models still exhibit significant room for improvement. These findings underscore the unique challenges posed by MDAR and its value as a benchmark for advancing audio reasoning research. Overall, our contributions are three-fold:
@@
    We introduce MDAR, a large-scale benchmark focusing on evaluating multi-scene and dynamic audio reasoning across five categories, spanning three question types, and for the first time proposing multiple-choice questions with multiple audios.
@@
    We construct a high-quality data-construction workflow and provide carefully curated audio clips paired with human-annotated questions and answers to systematically evaluate both perceptual and high-level reasoning abilities.
@@
    We conduct a comprehensive evaluation of state-of-the-art models, revealing the significant challenges posed by MDAR and highlighting key areas for improvement in next-generation audio reasoning agents.
@@ 2.1 Audio-Language Models
Recent advances in large language models (LLMs) and cross-modal learning have driven substantial progress in audio-language research. Typical approaches jointly leverage speech, music, and general audio data, followed by instruction tuning to endow models with comprehension and instruction-following abilities. Open-source systems such as Qwen-Audio-Chat (Chu et al., 2023), Qwen2-Audio-Instruct (Chu et al., 2024), Audio Flamingo (Ghosh et al., 2025; Goel et al., 2025), SALMONN (Tang et al., 2024), and DeSTA2.5-Audio (Lu et al., 2025) report competitive results. More recent efforts, including Audio-CoT (Ma et al., 2025a), Audio-Reasoner (Xie et al., 2025), and R1-AQA (Li et al., 2025), focus on multi-step reasoning and challenging audio QA. Among proprietary systems, GPT-4o-Audio achieves the highest overall performance, while other commercial models such as Kimi-Audio (KimiTeam et al., 2025) and MiDashengLM (Dinkel et al., 2025) are also evaluated. Fully multimodal models such as Omni-R1 (Zhong et al., 2025) and Qwen2.5-Omni (Xu et al., 2025a) advance unified audio-text reasoning. To provide a holistic comparison, our study additionally benchmarks cascaded audio-language pipelines, enabling a systematic evaluation of state-of-the-art approaches.
@@ 2.2 Audio Understanding and Reasoning Benchmarks
Prior works have explored audio question answering and compositional reasoning across speech, music, and environmental sounds. Early efforts include Lipping et al. (2022), a crowdsourced dataset with yes/no and single-word answers, and Ghosh et al. (2024b), which targets order- and attribute-level compositional reasoning with composition-aware fine-tuning for CLAP (Wu et al., 2023). In music, Melechovský et al. (2024) provides a controllable text-to-music system with theory-informed captions, and Weck et al. (2024) offers a human-validated multiple-choice benchmark probing music knowledge. For speech, Zhao et al. (2024) curates large-scale spoken QA in free-form and multiple-choice formats, while Huang et al. (2023) establishes a collaborative instruction-tuning benchmark across diverse speech tasks. Broader evaluations include Wang et al. (2025a), covering speech, scenes, and paralinguistics under instruction following, and Yang et al. (2024), which assesses generative comprehension and chat-based interaction over speech, sounds, and music. More recent reasoning-focused suites such as Sakshi et al. (2025), Ma et al. (2025b), and Wang et al. (2025b) emphasize multi-step perception and domain knowledge, revealing persistent gaps in multimodal integration and deep audio reasoning. However, these benchmarks mainly target single-scene or static audio scenarios and do not evaluate dynamic, multi-scene audio reasoning, which MDAR aims to address. MDAR introduces dynamic mixed-modal audio scenes, providing a more realistic and challenging benchmark compared to MMAU and MMAR.
3 MDAR benchmark
@@ 3.1 Overview of Benchmark
Figure 2: Examples from MDAR exquisitely showcase the diversity of the five complex dynamic reasoning tasks and also reveal the design of our newly proposed multiple-choice questions. These tasks not only demonstrate the depth of MDAR in multidimensional capability assessment but also reflect its high standards of challenge.
Statistics Number
Total Questions 1500
Task 5
MDAR-main Sub task 9
Average question length 30.68
Average option length 18.69
Average audio length 25.11
Total Questions 500
Task 5
MDAR-open Average question length 41.47
Average reference length 175.53
Average audio length 25.14
Total Questions 825
Task 3
Average question length 32.80
MDAR-multi Average answer length 47.19
Average audio length 69.90
Average answer number 2.82
Figure 3: (a) Distribution of all types of tasks included in the three benchmark tests illustrates the diversity and complexity of MDAR tasks. (b) Statistical information of the three benchmark tests indicates the complexity of the benchmarks through the longer duration of the audio and the rich semantic information.
@@
Overview. MDAR is a benchmark test specifically designed to evaluate the reasoning capabilities of audio-language models in complex and dynamic scenarios. It encompasses 3,000 questions involving high-quality complex reasoning tasks of various types. Figure 2 illustrates examples of different categories and types of questions within this benchmark test. Each example consists of a meticulously designed question and a reference answer. The questions are semi-automatically generated by LLMs and manually annotated by experts and have undergone multiple rounds of screening to ensure their high quality. This benchmark test poses extremely high demands on the complex dynamic reasoning abilities of models, presenting a significant challenge.
Categories and Subcategories. Single audio signals or static scenes are insufficient for a comprehensive evaluation of speech agents’ reasoning ability. We design three task types: single-choice questions for precise reasoning on complex audio, multiple-choice questions for cross-audio integrative analysis, and open-ended questions for deep reasoning and generation in dynamic scenarios.
@@
MDAR-main consists of 1,500 carefully designed single-choice questions, organized into five categories: scene understanding, social relationships and social reasoning, event reasoning, temporal reasoning, and anomaly detection and safety. Detailed descriptions and evaluation focuses for each category are provided in the Appendix B.
@@
MDAR-open anchors its question bank in real-world events and scenarios. The traditional closed-set multiple-choice evaluation method can no longer adequately reflect the core capabilities required by AI agents in real-world open environments. The uncertainty inherent in open-ended questions poses an even greater challenge to the agents’ authentic response and judgment abilities. Task categories are the same as MDAR-main.
@@
MDAR-multi is the first benchmark to incorporate multi-audio, multiple-choice questions. By embedding real-world events and semantic ambiguities into options, it challenges models to make interpretable decisions even under mishearing, omission, or bias. High-order reasoning tasks are grouped into three dimensions: plot development reasoning, Continuous scene reasoning, and multi-character interaction reasoning, emphasizing knowledge integration and decision-making in real-world contexts, thus providing a finer-grained and more realistic challenge.
@@
Question Distribution and Difficulty. Figure 3 shows the distribution of questions across different categories and subcategories and summarizes the key statistics of MDAR. All questions require a comprehensive combination of perception, understanding, and reasoning. The average lengths of questions and answers are 30.68 words and 18.69 words, respectively. The average length of the audio clips is 25.11 seconds,longer than the previous benchmark MMAU which is about 10 seconds and MMAR, which is about 19.94 seconds. Moreover, the three types of questions in MDAR go beyond the single multiple-choice question type in MMAU and MMAR, offering a more complex and comprehensive measure of audio reasoning capabilities.
3.2 Data construction pipeline
As illustrated in the Figure 4, we constructed a high-quality data-construction workflow, which is divided into three main steps. The details of prompts are provided in the Appendix D.
Data Preparation. MDAR uses Chinese movies as its metadata source, leveraging multi-threaded narratives and high-production films to construct complex dynamic scenes characterized by openness, high entropy, long temporal dependencies, and strong causality. The metadata set contains 500 films covering a variety of genres and themes (Details are provided in Appendix C). We implemented three sub-processes: (i) Segment Sampling. Randomly cropped 20-40-second segments from target films to fit the audio input limitations of existing LALMs while ensuring sufficient length for events and complex information. (ii) Speaker Diarization. Using the open-source tool Pyannote to label speaker information in audio, ensuring that multiple speech segments have clear scene events and rich information. (iii) Global Segment Clustering. Clustered audio segments with the same speaker to form the audio pairs needed for multiple-choice questions.
Audio Pipeline. Based on the above corpus, we design a five-step pipeline: (i) Information Preparation. We apply FunASR for speech recognition and use Gemini-2.5-pro and Qwen2.5-VL to generate multimodal descriptions, laying the foundation for question construction. (ii) Framework Construction. Domain experts are invited to design question types and categories, and to select complex, dynamic scenarios that sufficiently challenge large audio language models (LALMs). (iii) Content Generation. Gemini-2.5-pro is used to automatically generate QA pairs, with task-specific prompts tailored for each task type to encourage diversity. For multi-audio QA, special attention is paid to combining independent audio clips with shared contextual information. (iv) Human-designed screening. Experts refine and annotate the QA pairs, producing complete metadata including identifiers, timestamps, questions, and answers. Rigorous quality control criteria are applied to remove invalid or irrelevant QA pairs, ensuring that all questions focus on audio content. More screening criteria are provided in the Appendix E. (v) Distractor generation. Large language models are used to generate distractor options. Each single-choice question contains four options, while in multiple-choice questions, distractors are evenly distributed between information from independent and shared audio clips. The final benchmark is output in a standardized JSON format.
Data Quality Assurance. We implement multiple traceable steps to ensure high-quality data: (i) Expert Annotation Filtering. Manually screening each sample to remove invalid, low-quality, harmful, or irrelevant ones to enhance data purity. Cross-checking by multiple annotators is employed to avoid subjective bias. (ii) Re-annotation. Manually correcting original annotation errors and standardizing annotation criteria to improve label accuracy. (iii) Secondary Review. Conducting a second screening according to data screening standards to ensure no omissions and no systematic biases. (iv) Category Balancing. Ensuring a balanced distribution of question categories to improve the quality of the test benchmark. If category imbalance occurs, returning to the audio process for category completion and iterating until balance is achieved.
3.3 Evaluation metrics
Table 1: Results of models on the MDAR-main benchmark for both audio language models and cascaded models are presented for each category. The best results are highlighted in bold and the second-best is underlined. The results cover five task categories: SR (Social Relationships and Social Reasoning), ER (Event Reasoning), SU (Scene Understanding), AD (Anomaly Detection and Safety), and TR (Temporal Reasoning).
@@ 3.3 Evaluation metrics
Qwen2-Audio-Instruct  (Chu et al., 2024) 7B LALMs 38.07 31.13 23.12 35.06 27.14 33.60
Qwen-Audio-Chat (Chu et al., 2023) 8.4B LALMs 17.93 20.34 10.98 15.52 22.86 17.73
@@
Kimi-Audio-Instruct (KimiTeam et al., 2025) 7B LALMs 14.52 18.38 13.29 20.69 25.71 16.67
Omni-R1 (Zhong et al., 2025) 7B OLMs 44.89 45.83 45.66 60.34 38.57 46.73
R1-AQA (Li et al., 2025) 7B LALMs 39.11 41.18 27.17 35.06 38.57 37.80
SALAMONN (Tang et al., 2024) 7B LALMs 36.45 29.76 37.55 35.29 37.89 34.75
Audio-Reasoner (Xie et al., 2025) 8.4B LALMs 45.93 42.65 43.35 37.36 47.14 43.80
@@
In this section, we introduce the evaluation methods for MDAR. Since MDAR has different question types, in order to maintain fairness and consistency, we adopt different evaluation methods.
MDAR-main. We use accuracy as the evaluation metric. The model is fed with audio, questions, and text instructions to generate answers. Since audio-language models (LALMs) are more adept at producing open-ended responses, we adopt the same response processing pipeline as MMAU. We use robust regular expressions to match the answer strings with the given options and then calculate the accuracy. We found that different prompts can affect the final results. In this evaluation, we used the prompt that yielded the highest average model performance as the final prompt. The results of option distribution and instruction bias are provided in the Appendix I.
MDAR-open. We employ a state-of-the-art LLM as an automated evaluator. The evaluation consists of four core components: prompt, question, answer, and reference answer. During the scoring process, the evaluator rates the model’s answer on a scale of 0 to 10 based on the given scoring prompt. These prompts consider the usefulness, relevance, accuracy, and comprehensiveness of the answer. Detailed prompt design and scoring examples are provided in the Appendix F.2. We reduce the randomness of scoring by conducting multiple evaluations and swapping the positions of the answers.
MDAR-multi. Inspired by SATA-BENCH (Xu et al., 2025b), we use the following four evaluation metrics to measure model performance. For answer set extraction, we still use robust regular expressions to ensure accuracy. (i) Exact Match (EM): For each sample, if the model’s predicted answer set is exactly the same as the correct answer set, the accuracy is counted as 1; otherwise, it is counted as 0. The final EM is calculated as the average accuracy. (ii) Jaccard Index (JI): For each sample, we calculate the intersection of the predicted and true answers divided by their union, and then take the average of all samples to measure the degree of overlap. (iii) Mean Average Precision (Precision): For each sample, we calculate the proportion of correctly predicted answers out of all predicted answers, and then take the average of all samples. (iv) Mean Average Recall (Recall): For each sample, we calculate the proportion of correctly predicted answers out of all true answers, and then take the average of all samples. The specific formulas, are provided in the Appendix F.1.
4 Experiments
Table 2: Scores of models on the MDAR-open benchmark for audio language models are presented for each category. The best results are highlighted in bold and the second-best is underlined. The five categories are represented in the same way as in Table $3.$

### v2

# CMDAR: A Chinese Multi-scene Dynamic Audio Reasoning Benchmark with Diverse Challenges

@@ Abstract
The ability to reason from audio, including speech, environmental sounds, and music, is essential for AI agents to interact effectively in real-world scenarios. Existing benchmarks mainly focus on static or single-scene settings and English audio data and do not fully capture scenarios where multiple speakers, unfolding events, and heterogeneous audio sources interact. To address these challenges, we introduce CMDAR, a chinese benchmark for evaluating models on complex, multi-scene, and dynamically evolving audio reasoning tasks. CMDAR comprises 3,000 carefully curated question–answer pairs linked to diverse audio clips, covering five categories of complex reasoning and spanning three question types. We benchmark 26 state-of-the-art audio language models on CMDAR and observe that they exhibit limitations in complex reasoning tasks. In CMDAR-main, Qwen2.5-Omni (open-source) achieves 76.67% accuracy, whereas GPT-4o Audio (closed-source) reaches 68.47%. However, GPT-4o Audio substantially outperforms Qwen2.5-Omni on the more challenging multiple-choice with multiple audios and open-ended tasks. And we provide detail analysis corresponding suggestions for the future development of large audio language models (LALMs).
: https://huggingface.co/datasets/…….
: https://github.com/luckyerr/CMDAR
CMDAR: A Chinese Multi-scene Dynamic Audio Reasoning Benchmark with Diverse Challenges
Hui Li1thanks:   Equal Contribution., Changhao Jiang1∗, Hongyu Wang2∗, Ming Zhang1, Jiajun Sun1, Zhixiong Yang1, Yifei Cao1, Shihan Dou1, Xiaoran Fan1, Baoyu Fan3, Tao Ji122footnotemark: 2, Tao Gui122footnotemark: 2, Qi Zhang1thanks:   Corresponding Author. , Xuanjing Huang122footnotemark: 2 1 College of Computer Science and Artificial Intelligence, Fudan University 2 Shanghai Jiao Tong University 3 IEIT Systems Co Ltd hui_li25@fudan.edu.cn
@@ 1 Introduction
Figure 1: The differences between the existing benchmark and CMDAR. Tasks in CMDAR are essential skills for AI Agent.
@@
So far, various benchmarks have been proposed. AudioBench (Wang et al., 2025a) and AIR-Bench (Yang et al., 2024) are benchmark designed for large audio language models, aiming to comprehensively evaluate the understanding and reasoning ability of these models on audio content, but only focusing on shallow reasoning ability. MMAU (Sakshi et al., 2025) and MMAR focus more on the evaluation of deep reasoning ability of audio, but they pay more attention to perception samples, including counting and classification perception tasks, mainly for single scene and single audio problems, and do not include the evaluation of open-ended tasks, resulting in insufficient evaluation. Recently, MMAU-pro (Kumar et al., 2025) addressed the evaluation problem of multi-audio, but all the above evaluations only focus on English data, and do not focus on the investigation of Chinese audio reasoning ability and instruction following ability. To address the limitations of current audio reasoning benchmarks, we propose CMDAR, a Chinese comprehensive test benchmark aimed at evaluating the complex logical reasoning ability of LALMs in dynamic scenarios. CMDAR has 3000 high-quality questions from real scenarios, with multiple types of reasoning tasks (including continuous scenarios, causal scenarios, and multi scenario tasks) and comprehensive question types. All questions and reference answers are strictly reviewed by experts. The differences between the existing benchmark and CMDAR are summarized in Figure 1. In addition, we also propose a high-quality data construction process and a comprehensive and effective evaluation method.
We benchmark 26 state-of-the-art audio language models on CMDAR and observe that they exhibit limitations in complex reasoning tasks. In the CMDAR-main, Qwen2.5-Omni (open-source) achieves 76.67% accuracy, whereas GPT-4o Audio (closed-source) reaches 68.47%. Among different tasks of single-choice questions, Qwen2.5-Omni (Xu et al., 2025a) performs worst on temporal reasoning (71.43%), while GPT-4o Audio struggles most on scene reasoning (61.27%). However, GPT-4o Audio substantially outperforms Qwen2.5-Omni on the more challenging multiple-choice with multiple audios and open-ended tasks. Across all three question types, no model surpasses 80% performance. These results indicate that even the most powerful models currently still have significant room for improvement, and finally we provide corresponding suggestions for the future development of LALMs. These findings underscore the unique challenges posed by CMDAR and its value as a benchmark for advancing audio reasoning research. Overall, our contributions are three-fold:
@@
    We introduce CMDAR, a first large-scale Chinese benchmark focusing on evaluating multi-scene and dynamic audio reasoning across five categories, spanning three question types.
@@
    We construct a high-quality data construction workflow and provide curated audio clips paired with human re-annotated questions and answers to systematically evaluate both perceptual and high-level reasoning abilities.
@@
    We conduct a comprehensive evaluation of state-of-the-art models, revealing the significant challenges posed by CMDAR and give suggestions of highlighting key areas for improvement in next-generation audio reasoning agents.
Figure 2: Examples from CMDAR showcase the diversity of complex dynamic reasoning tasks and also reveal the design of our newly proposed multiple-choice questions. These tasks not only demonstrate the depth of CMDAR in multidimensional capability assessment but also reflect its high standards of challenge.
@@ 2.1 Audio-Language Models
Recent advances in large language models (LLMs) and cross-modal learning have driven substantial progress in audio-language research. Typical approaches jointly leverage speech, music, and general audio data, followed by instruction tuning to endow models with comprehension and instruction-following abilities. Open-source systems such as Qwen-Audio-Chat (Chu et al., 2023), Qwen2-Audio-Instruct (Chu et al., 2024), Audio Flamingo (Ghosh et al., 2025; Goel et al., 2025), SALMONN (Tang et al., 2024), and DeSTA2.5-Audio (Lu et al., 2025) report competitive results. More recent efforts, including Audio-CoT (Ma et al., 2025a), Audio-Reasoner (Xie et al., 2025), and R1-AQA (Li et al., 2025), focus on multi-step reasoning and challenging audio QA. Among proprietary systems, GPT-4o-Audio achieves the highest overall performance, while other commercial models such as Kimi-Audio (KimiTeam et al., 2025) and MiDashengLM (Dinkel et al., 2025) are also evaluated. Fully multimodal models such as Omni-R1 (Zhong et al., 2025) and Qwen2.5-Omni (Xu et al., 2025a) advance audio-text reasoning. We also benchmark cascaded audio-language models for a holistic comparison.
@@ 2.2 Audio Understanding and Reasoning Benchmarks
Prior works have explored audio question answering and compositional reasoning across speech, music, and environmental sounds. Early efforts include Lipping et al. (2022), a crowdsourced dataset with yes/no and single-word answers, and Ghosh et al. (2024b), which targets order- and attribute-level compositional reasoning with composition-aware fine-tuning for CLAP (Wu et al., 2023). In music, Melechovský et al. (2024) provides a controllable text-to-music system with theory-informed captions, and Weck et al. (2024) offers a human-validated multiple-choice benchmark probing music knowledge. For speech, Zhao et al. (2024) curates large-scale spoken QA in free-form and multiple-choice formats, while Huang et al. (2023) establishes a collaborative instruction-tuning benchmark across diverse speech tasks. Broader evaluations include Wang et al. (2025a), covering speech, scenes, and paralinguistics under instruction following, and Yang et al. (2024), which assesses generative comprehension and chat-based interaction over speech, sounds, and music. More recent reasoning-focused suites such as Sakshi et al. (2025), Ma et al. (2025b), and Wang et al. (2025b) emphasize multi-step perception and domain knowledge, revealing persistent gaps in multimodal integration and deep audio reasoning.
3 CMDAR benchmark
@@ 3.1 Overview of Benchmark
CMDAR is a benchmark test specifically designed to evaluate the reasoning capabilities of audio-language models in complex and dynamic scenarios. It encompasses 3,000 questions involving high-quality complex reasoning tasks of various types. Figure 2 illustrates examples of different categories and types of questions within this benchmark test. Each example consists of a meticulously designed question and a reference answer. The questions are semi-automatically generated by LLMs and manually annotated by experts and have undergone multiple rounds of screening to ensure their high quality. This benchmark test poses extremely high demands on the complex dynamic reasoning abilities of models, presenting a significant challenge.
@@
CMDAR-main consists of 1,500 carefully designed single-choice questions for precise reasoning on complex audio, organized into five categories: scene understanding, social relationships and social reasoning, event reasoning, temporal reasoning, and anomaly detection and safety. Detailed descriptions and evaluation focuses for each category are provided in the Appendix A.
@@
CMDAR-open anchors its question bank in real-world events and scenarios. The traditional closed-set multiple-choice evaluation method can no longer adequately reflect the core capabilities required by AI agents in open real-world environments. The uncertainty inherent in open-ended questions poses greater challenge to the agents’ authentic response and judgment abilities. Task categories are the same as CMDAR-main.
@@
CMDAR-multi is the benchmark to incorporate multi-audio, multiple-choice questions. By embedding real-world events and semantic ambiguities into options, it challenges models to make interpretable decisions even under mishearing, omission, or bias. Detailed descriptions and evaluation focuses for each category are provided in the Appendix A.
@@
Question Distribution and Difficulty. Figure 5 shows the distribution of questions across different categories and subcategories and summarizes the key statistics of CMDAR. The average lengths of questions and answers are 30.68 words and 18.69 words. The average length of the audio clips is 25.11 seconds,longer than the previous benchmark MMAU which is about 10 seconds and MMAR, which is about 19.94 seconds. Moreover, the three types of questions in CMDAR go beyond only the single-choice question type in MMAU and MMAR, offering a more complex and comprehensive measure of audio reasoning capabilities.
[…]
@@
[…]
@@
[…]
@@
3.2 Data construction pipeline
As illustrated in Figure 3, we propose a high-quality data-construction workflow, which is divided into three main steps. The details of prompts are provided in the Appendix D.
Data Preparation. CMDAR uses Chinese films as metadata source, covering various genres and themes (Details are provided in Appendix B). Leveraging multi-narrative and high-production-value films to construct complex and dynamic scenes, we randomly extracted 20-40 seconds clips from the target films to accommodate the existing limitations of audio input, while ensuring the clip length is sufficient to contain events and complex information. Then we use PyAnnote toolkit [^1] https://github.com/pyannote/pyannote-audio to segment the clips and perform global clustering of audio clips with the same speaker to form the audio pairs required for multiple-choice questions.
Audio Pipeline. Based on the above corpus, we design a three-step pipeline: i) Information Preparation. We apply FunASR [^2] https://github.com/modelscope/FunASR for speech recognition and use Gemini-2.5-pro and Qwen2.5-VL to generate multimodal descriptions, laying the foundation for question construction. ii) Framework Construction. Domain experts are invited to design question types and categories, and to select complex, dynamic scenarios that sufficiently challenge LALMs. iii) Content and Distractor Option Generation. We use Gemini-2.5-pro to automatically generate question-answer pairs and distractor options. For multi-audio Q&A, we focus on combining independent audio segments with shared contextual information, while distractor options for multiple-choice questions were evenly distributed among the information from independent audio segments and shared audio segments.
Data Quality Assurance. We implement multiple traceable steps to ensure high-quality data: i) Expert Annotation Filtering. Experts refine and annotate the Q&A pairs, generating complete metadata including identifiers, timestamps, questions, and answers. We use strict standards to manually screen for invalid, low-quality, harmful, or irrelevant samples, ensuring that all issues revolve around the audio content. More screening criteria are provided in the Appendix E. ii) Secondary Review. Conducting a second screening according to data screening standards to ensure no omissions and no systematic biases. iii) Category Balancing. Ensuring a balanced distribution of question categories to improve the quality of the test benchmark. If not, returning to Step2 until balance is achieved.
Table 2: Scores of audio language models on the CMDAR-open benchmark are presented for each category. The best results are highlighted in bold and the second-best is underlined. The five categories are represented in the same way as in Table 1.